OpenStudy (helpmeee):

Giselle graphs the function f(x) = x2. Robin graphs the function g(x) = –x2. How does Robin’s graph relate to Giselle’s? Robin’s graph is a reflection of Giselle’s graph over the x-axis. Robin’s graph is a reflection of Giselle’s graph over the y-axis. Robin’s graph is a translation of Giselle’s graph 1 unit down. Robin’s graph is a translation of Giselle’s graph 1 unit left.

OpenStudy (mathmale):

First, some housekeeping: Exponentiation is expressed using the symbol " ^ " so your x2 should be written as x^2, "the square of x." Giselle graphs the function f(x) = x^2. Robin graphs the function g(x) = –x^2. How does Robin’s graph relate to Giselle’s? Please graph both y=x^2 and y=-x^2 on the same set of axes. Use the Draw utility, below.

OpenStudy (helpmeee):

that option does not work on my computer

OpenStudy (helpmeee):

@mathmale

OpenStudy (helpmeee):

@mathmate

OpenStudy (mathmale):

Please find some other way of sharing your work. You could, for example, draw the graph of y=x^2 on paper, photograph it with your cell phone, and then upload the image to OpenStudy.

OpenStudy (mathmate):

@helpMEE Please continue with @mathmale

OpenStudy (helpmeee):

thats so much stuff i only have 15mins left in my exam

OpenStudy (helpmeee):

can you trust me

OpenStudy (mathmate):

@helpMEE Students are NOT ALLOWED to use OS for time-limited evaluations of any kind, exams, quizzes, etc.

OpenStudy (helpmeee):

its not thats serious